# CPT: Competence-progressive Training Strategy for Few-shot Node Classification

### v1

@@ Abstract
Graph Neural Networks (GNNs) have made significant advancements in node classification, but their success relies on sufficient labeled nodes per class in the training data. Real-world graph data often exhibits a long-tail distribution with sparse labels, emphasizing the importance of GNNs’ ability in few-shot node classification, which entails categorizing nodes with limited data. Traditional episodic meta-learning approaches have shown promise in this domain, but they face an inherent limitation: it might lead the model to converge to suboptimal solutions because of random and uniform task assignment, ignoring task difficulty levels. This could lead the meta-learner to face complex tasks too soon, hindering proper learning. Ideally, the meta-learner should start with simple concepts and advance to more complex ones, like human learning. So, we introduce CPT, a novel two-stage curriculum learning method that aligns task difficulty with the meta-learner’s progressive competence, enhancing overall performance. Specifically, in CPT’s initial stage, the focus is on simpler tasks, fostering foundational skills for engaging with complex tasks later. Importantly, the second stage dynamically adjusts task difficulty based on the meta-learner’s growing competence, aiming for optimal knowledge acquisition. Extensive experiments on popular node classification datasets demonstrate significant improvements of our strategy over existing methods.
@@ 1 Introduction
In recent years, there has been a lot of research focused on node classification, which is about predicting labels for unlabeled nodes in a graph. This task has many practical applications in real-world situations Yan et al. (2021); Zhang et al. (2020). For example, in bioinformatics Szklarczyk et al. (2019), predicting chemical properties for proteins in a protein network is an important problem that can be tackled using node classification. State-of-the-art methods for node classification often use Graph Neural Networks (GNNs)Wu et al. (2021); Xu et al. (2018) in a semi-supervised manner Kipf and Welling (2016). GNNs aim to learn representations for each node by considering information from its neighbors. These representations are then used for classification. However, these methods usually require a sufficient number of labeled nodes for all classes to achieve good results Fan et al. (2019). In practice, we often have many labeled nodes for some classes, but only a few for others. We refer to classes with many labels as base classes and classes with few labels as novel classes Wang et al. (2022). Since novel classes are common in real-world graphs, it’s important for GNNs to be capable of classifying nodes even with limited labeled nodes, known as the few-shot node classification problem.
@@
Most existing studies on the few-shot node classification utilize the episodic meta-learning strategy Wang et al. (2022, 2020); Liu et al. (2021a). In this paradigm, the meta-learner learns from base classes through a series of meta-training tasks and evaluates the model using meta-test tasks derived from novel classes. The aim of this scheme is to allow the meta-learner to extract meta-knowledge from base classes, enabling it to classify using just a few labeled nodes from the novel classes. Despite the success of existing methods using the episodic meta-learning scheme, the episodic training strategy would introduce some issues. First and foremost, a fundamental limitation of episodic training methods is their random selection of tasks from base classes, neglecting the varied difficulty levels of these tasks. This differs significantly from the natural human learning process, where we typically begin with basic concepts and progressively advance to more complex ones. As a result, in such a training model, there’s a risk that the meta-learner might be exposed to challenging tasks too early and simpler tasks later, potentially leading to less effective knowledge acquisition. As shown in Figure 1, when the model is first initialized (akin to a toddler phase), it is not yet capable of handling difficult tasks, but it can acquire knowledge from easy tasks. At this stage, easy tasks are informative and beneficial for the model. After sufficient training, the model develops intelligence to tackle more complex tasks, resembling a teenager phase. By then, easy tasks become uninformative for it, and it is the hard tasks that truly contribute to enhancing its intelligence. Inspired by the human learning process, the idea of learning from easy to hard has been effectively implemented in curriculum learning Hacohen and Weinshall (2019) and self-paced learning Jiang et al. (2014). In these frameworks, individual samples are given different levels of importance (i.e., weights) to enhance generalization. Although the method of using curriculum learning to address problems in few-shot tasks has been studied Sun et al. (2019); Liu et al. (2020), it is not applicable to graph data. Graph data has its own complex relational dependency patterns, and existing work cannot be directly transferred over.
@@
  • Problem. We explore the limitations of the episodic meta-learning scheme and discuss its associated issues in few-shot node classification. Specifically, the limitation of the episodic meta-learning scheme is that it samples tasks randomly and uniformly, which can lead the meta-learner to converge to a suboptimal point and thus make further optimization more challenging.
@@
  • Experiments. We conduct experiments on four benchmark node classification datasets under the few-shot setting and demonstrate significant improvements of our strategy over existing methods.
2 Related Work
2.1 Few-shot Learning on Graphs
2.2 Curriculum Learning

### v3

@@ Abstract
Traditional episodic meta-learning methods have made significant progress in solving the problem of few-shot node classification on graph data. However, they suffer from an inherent limitation: the high randomness in task assignment often traps the model in suboptimal solutions. This randomness leads the model to encounter complex tasks too early during training, when it still lacks the necessary capability to handle these tasks effectively and acquire useful information. This issue hampers the model’s learning efficiency and significantly impacts the overall classification performance, leading to results that fall well below expectations in data-scarce scenarios. Inspired by human learning, we propose a model that starts with simple concepts and gradually advances to more complex tasks, which facilitates more effective learning. Therefore, we introduce CPT, a novel curriculum learning strategy designed for few-shot node classification. This approach progressively aligns the meta-learner’s capabilities with both node and task difficulty, enhancing the learning efficiency for challenging tasks and further improving overall performance. Specifically, CPT consists of two stages: the first stage begins with basic tasks, aligning the difficulty of nodes with the model’s capability to ensure that the model has a solid foundational ability when facing more complex tasks. The second stage focuses on aligning the difficulty of meta-tasks with the model’s capability and dynamically adjusting task difficulty based on the model’s growing capability to achieve optimal knowledge acquisition. Extensive experiments on widely used node classification datasets demonstrate that our method significantly outperforms existing approaches.
@@ 1 Introduction
In recent years, there has been a lot of research focused on node classification, which is about predicting labels for unlabeled nodes in a graph. This task has many practical applications in real-world situations. For example, it can help us identify molecules that play a crucial role in protein structures Szklarczyk et al. (2019), recognize important influencer in social networks Yuan et al. (2013); Zhang et al. (2020), or infer users’ demographic information based on behavior patterns in recommendation systems Yan et al. (2021). Although Graph Neural Networks (GNNs) have been successful in node classification tasks Wu et al. (2021); Xu et al. (2018), their performance heavily relies on the number of labeled nodes in each class Fan et al. (2019). In real-world graphs, we often have a large number of labeled nodes in some classes, while other classes have only a limited number of labeled nodes. We refer to classes with a large number of labeled samples as base classes, and those with a limited number of labels as novel classes Wang et al. (2022). For example, in disease diagnosis, common diseases with abundant samples can be considered base classes, while newly discovered diseases with limited samples are regarded as novel classes. Due to the widespread existence of novel classes in real-world graphs, it is important for GNNs to be able to perform node classification with a limited amount of labeled nodes, a task known as few-shot node classification.
@@
Unlike the traditional node classification problem, few-shot node classification has training data with only a small number of samples, which leads to the challenge of insufficient knowledge. The current mainstream learning paradigm addresses this issue by constructing meta-tasks through meta-learning to transfer knowledge Wang et al. (2022, 2020); Liu et al. (2021a). In this paradigm, the meta-learner learns from a series of meta-training tasks derived from the base classes and evaluates the model using meta-testing tasks derived from the novel classes. The goal of this approach is to enable the meta-learner to extract meta-knowledge from the base classes, allowing it to classify novel classes with only a few labeled nodes. Although this meta-learning paradigm has achieved success in the few-shot node classification task, it has inherent flaws in its context-based training strategy. This strategy randomly selects meta-tasks for learning, which may lead to certain issues. In node classification tasks, one of the primary factors affecting overall performance is learning from difficult nodes, and a similar situation exists in few-shot graph learning, as shown in Figure 2. Since meta-tasks are composed of nodes in meta-learning, there are also differences in task difficulty. The context-based training strategy randomly selects tasks from the base classes for learning, ignoring the differences in task difficulty. This may cause a mismatch between the meta-learner’s capability and the task’s difficulty, leading to suboptimal performance when learning from difficult tasks and thereby affecting overall results.
This idea is inspired by the human learning process, where humans typically start learning from basic concepts and gradually progress to more complex content, as harder tasks require knowledge from simpler ones as a foundation. This strategy of progressively learning from easy to hard tasks has been effectively applied in curriculum learning Hacohen and Weinshall (2019) and adaptive learning Jiang et al. (2014). In these frameworks, individual samples are assigned different levels of importance (i.e., weights) to enhance the model’s generalization ability. In other words, randomly selecting tasks may expose the meta-learner to high-difficulty tasks too early, resulting in ineffective acquisition of useful information from those challenging tasks. As shown in the Figure 1, at the initial stage (analogous to an infant stage), the model lacks the capability to handle difficult tasks but can acquire knowledge from simpler tasks. At this stage, simple tasks are rich in information and beneficial to the model. After sufficient training, the model’s intelligence level improves, resembling an adolescent stage, where learning from difficult tasks becomes more effective due to its foundational capabilities. While previous studies have explored the use of curriculum learning in node classification tasks Wei et al. (2023), their approaches are primarily focused on nodes and cannot be applied to meta-tasks, such as node-level oversampling or reparameterization. Although there are curriculum learning approaches designed for meta-tasks Sun et al. (2019); Liu et al. (2020), they focus on image and text data, without considering the complex relational dependencies in graph data. Therefore, new solutions are needed to address the challenges of few-shot classification on graph data.
@@
  • Problem. We explore the limitations of context-based meta-learning approaches and discuss their relevant issues in few-shot node classification. Specifically, the limitation of these approaches lies in their random and uniform sampling of tasks, which may result in insufficient training of the meta-learner on difficult tasks, leading to suboptimal model performance.
@@
  • Experiments. We conducted experiments on four benchmark node classification datasets under few-shot settings, and the results demonstrate that our strategy significantly improves performance over existing methods.
@@ 4.7 Ablation Study
5 Related Work
5.1 Few-shot Learning on Graphs
5.2 Curriculum Learning